Two angles are complementary. The sum of the measure of the first angle and one- fourth the second angle is 82.5 degree. Find the measures of the angles
Answer by mananth(16946)

Two angles are complementary. The sum of the measure of the first angle and one- fourth the second angle is 82.5 degree. Find the measures of the angles
...
let one angle =x
the pther angle = 90-x
..
x+1/4*(90-x)=82.5
(4x+90-x)/4=82.5
multiply by 4
3x+90=4*82.5
3x+90=330
-90
3x=330-90
3x=240
/3
x=80 degrees
other angle = 10 degrees
..
m.ananth@hotmail.ca